Suppose MPC=.75. What is the change in G if you want Y to increase by $2,000?
Answer #1

If the MPC in the market is at 0.75 then the multiplier in the market will be at 1/ 1-MPC = 1/ 1-0.75 = 4.

For increasing the Y by 2000 then the government has to increase the output by 2000/4 = $500.